### v2

# Re-parameterized Low-rank Prompt: Generalize a Vision-Language Model within 0.5K Parameters

Abstract
With the development of large pre-trained vision-language models, how to effectively transfer the knowledge of such foundational models to downstream tasks becomes a hot topic, especially in a data-deficient scenario. Recently, prompt tuning has become a popular solution. When adapting the vision-language models, researchers freeze the parameters in the backbone and only design and tune the prompts. On the one hand, the delicate design of prompt tuning exhibits strong performance. On the other hand, complicated structures and update rules largely increase the computation and storage cost. Motivated by the observation that the evolution pattern of the generalization capability in visual-language models aligns harmoniously with the trend of rank variations in the prompt matrix during adaptation, we design a new type of prompt, Re-parameterized Low-rank Prompt (RLP), for both efficient and effective adaptation. Our method could largely reduce the number of tunable parameters and storage space, which is quite beneficial in resource-limited scenarios. Extensive experiments further demonstrate the superiority of RLP. In particular, RLP shows comparable or even stronger performance than the latest state-of-the-art methods with an extremely small number of parameters. On a series of tasks over 11 datasets, RLP significantly increases the average downstream accuracy of classic prompt tuning by up to 5.25% using merely 0.5K parameters.
1 Introduction
In recent years, vision-languages models [37, 17] have achieved tremendous success. Representative models like CLIP [37] are first pre-trained on a huge number of text-image pairs on the web to align textual and visual features, and then can be tuned and used for various downstream tasks.
Figure 1: Comparison between classic prompt tuning and Re-parameterized Low-rank Prompt (RLP) tuning. RLP introduces two little parameter matrices in shape $[n,r]$ and $[r,d]$ separately, and uses the product of them as an equivalent low-rank prompt in shape $[n,d]$.
However, traditional fine-tuning is not a good choice to adapt vision-language models. Simply fine-tuning all the parameters can easily cause the model to forget task-agnostic knowledge that is general and important, and the huge training and storage cost is also an intractable problem. Prompt engineering, a recent popular method in natural language processing and computer vision, provides an overwhelming solution for the adaptation of vision-language models. Prompts are virtual tokens that are manually appended or prepended to the semantic tokens. Such virtual tokens can be hand-crafted designed and frozen [41], and they can be soft and tunable as well [30, 27, 18] for better transferability on different tasks. Typical prompt tuning methods freeze all the other parameters in the model and merely tune the prompts during adaptation.
In the vision-language field, prompt-based methods show good performance as well. CLIP [37] uses a hand-crafted text template for better zero-shot learning. Bahng et al. [1] adopts prompt tuning merely on the image encoder, CoOp [54] uses tunable text prompts to replace the fixed template in CLIP, CoCoOp [55] utilizes image feature to instruct the optimization of the soft text prompt in CoOp.  [21, 24] simultaneously optimize image and text prompts and establish extra connections between isolated modalities.  [22, 47, 3, 53] integrate strong regularization modules or losses into prompt tuning to diminish the overfitting and catastrophic forgetting problem. On the one hand, such delicate designs of prompt tuning in recent methods exhibits strong performance. On the other hand, their complicated structures and update rules largely increase the computation and storage cost.
Motivated by the observation that the evolution pattern of the generalization capability in visual-language models aligns harmoniously with the trend of rank variations in the prompt matrix during adaptation as shown in Figs. 2 and 1, we propose Re-parameterized Low-rank Prompt (RLP) to reach effective and efficient adaptation for vision-language models. For more details about the motivation and our observation, please refer to Sec. 2 and we will provide a thorough analysis. In this paper, we have explored how to effectively adapt vision-language models under an extremely small number of parameters. By choosing a small rank $r$, we could exceedingly compress the parameters used in adaptation. To the best of our knowledge, we are the first the reach comparable or even better performance compared with the latest state-of-the-art tuning methods in the base-to-new generalization and domain generalization using such barely few parameters, i.e. 0.5K.
Additionally, though more complex and delicate, existing methods are still refined on a common fundamental basis, i.e. prompt tuning. Such a common basis guarantees that our RLP could be easily and smoothly integrated into most of the off-the-shelf methods besides individually applied.
In summary, we conclude our contributions as follows:
  • Motivated by the observation that the evolution pattern of the generalization capability in visual-language models aligns harmoniously with the trend of rank variations in the prompt matrix during adaptation, we propose to re-parameterize classic prompts into low-rank prompts for better accuracy and efficiency.
  • We propose a novel initialization method and integrate Dropout layer as a lightweight regularization module to further improve the performance of low-rank prompt tuning without introducing any extra parameters and inference cost.
  • We are the first to explore how to effectively adapt vision-language models using an extremely small number of parameters, i.e. 0.5K.
  • We conduct extensive experiments and show the fantastic effectiveness and efficiency of our proposed method. In base-to-new generalization, domain generalization, cross-dataset transfer and few-shot learning settings, RLP consistently gives very competitive results, surpassing many state-of-the-art tuning methods though we use fewer parameters.
2 Motivation for Low-Rank Prompts
In this section, we provide the motivation for our design of low-rank prompts. First of all, we do classic prompt tuning on a CLIP [37] model following CoOp [54] and save checkpoint for every epoch. The prompts are randomly initialized here for clear observation. Since we want to inspect the rank properties of the prompt matrix, we do singular value decomposition (SVD) for the prompt matrix $P\in\mathbb{R}^{n\times d}$.
$P=U\Sigma V^{T}$ (1)
where $\Sigma$ is the singular matrix and the values on its main diagonal denoted by $\sigma_{i}$ $(1\leq i\leq m)$ are singular values, which are non-negative. Here $m=min(n,d)$. The number of non-zero singular values is exactly the rank of the prompt matrix. In practice, even though all of the singular values are non-zero, the original matrix could be approximately considered a low-rank matrix as well if some singular values are far smaller than others. Therefore, we compute two values, ``std'' and ``scale'', to investigate the low-rank property of singular values in training time. At a certain point during training, we denote the largest singular value as $\sigma_{max}$ while the smallest singular value as $\sigma_{min}$. Since the SVD result is not unique, we do a scaling transformation first and then compute the standard deviation of the singular values.
$\forall 1\leq i\leq m,\sigma_{i}^{\prime}=\frac{\sigma_{i}}{\sigma_{min}}$ (2)
$\mbox{std}=\sqrt{\frac{1}{m}\sum_{i=1}^{m}{(\sigma_{i}^{\prime}-\frac{\sum_{j=1}^{n}\sigma_{j}^{\prime}}{m})^{2}}}$ (3)
We define ``scale'' as the ratio of the largest singular value and smallest singular value.
$\mbox{scale}=\frac{\sigma_{max}}{\sigma_{min}}$ (4)
Naturally, when the ``std'' is getting larger, the prompt matrix is getting closer to a low-rank form. The same holds true for the ``scale''.
Figure 2: Curves for the generalization ability (represented by ``new acc'' here) and the singular value distribution of prompt matrix. When ``new acc'' increases, ``std'' and ``scale'' also increase, i.e. the prompt matrix is getting closer to a low-rank form.
Table 1: Spearman correlation coefficients. Clearly, accuracy in the new classes is strongly correlated with std/scale of the singular values of the prompt matrix.
SUN397 DTD
std scale std scale
new acc 0.967 0.970 0.907 0.782
We show the result of each training epoch on base-to-new generalization setting in Fig. 2. The training curves are on two datasets, SUN397 [46] and DTD [4], respectively. Here we only show the test accuracy in the unseen new classes, which could measure the generalization ability of the model. For better visualization, we entirely scale up ``std'' to 8x and 6x on SUN397 and DTD respectively to match the value of ``scale'', so we could put them under the same axis in the figure. We also compute the Spearman correlation coefficient in Tab. 1, which measures the statistical dependence between the rankings of two variables. Seen from Figs. 2 and 1, the generalization ability of the vision-language model is highly correlated with ``std'' and ``scale'', e.g. the Spearman correlation coefficient of ``std'' and ``new acc'' is 0.967 on SUN397.
In summary, during training, the adaptation ability of the model represented by accuracy on the base classes is consistently improved as we have demonstrated in Fig. 5. The generalization ability of the model is improved in the first few epochs as well. However, as the training goes on, the generalization ability starts dropping, which indicates that the model suffers from overfitting and catastrophic forgetting problems. Seen from Fig. 2, the fluctuation trend of ``new acc'' is highly consistent with ``std'' and ``scale''. Naturally, we can infer that the low-rank form of the prompt matrix is beneficial to the generalization ability of the vision-language models. Motivated by such observation and inference, we propose Re-parameterized Low-rank Prompts (RLP) to get better generalization ability. Moreover, low-rank prompts naturally contain fewer parameters compared with classic ones, and thus we can also earn efficiency improvements.
3 Related Works
3.1 Vision-Language Models
Recently, large-scale vision-language models have shown very competitive performance in various tasks. Classic works [37, 17, 51, 48, 49] learn the multi-modal representation by a self-supervised manner on a large amount of image-text pairs. The representative work CLIP [37] is a milestone, which aligns the vision representation and language representation by contrastive learning and shows excellent performance. A well-trained vision-language model is a great treasure, which could largely facilitate the development of various related fields. There have been successful applications of such strong models on few-shot recognition [55, 54], detection [39, 32, 9, 50] and segmentation [26, 38, 6, 31]. For video data, there are also works on video classification [36] and video understanding [20].
3.2 Prompt Tuning
Originating from natural language processing, prompts are first introduced as a fixed template [41], e.g. a photo of a _, which is hand-crafted and fixed. Later, a series of methods [27, 25, 29, 42, 28, 19] are proposed to make such prompts tunable and be optimized during adaptation. Prompt tuning could adaptively narrow the gap between pre-trained representations and downstream tasks, significantly facilitating the fine-tuning process. Representative prompt tuning methods would add tunable virtual tokens, i.e. prompts, along with the semantic tokens as inputs of the model. All of the tokens are processed together to get text embeddings first and then sent to the feature encoder. Witnessing the success of prompting language models, researchers design prompts [18, 52] for visual models in a similar way. In vision-language field, there are several explorations as well. Bahng et al. [1] adopts prompt tuning merely on the image encoder. CoOp [54] uses tunable text prompts to replace the fixed template in CLIP [37]. CoCoOp [55] utilizes image feature to instruct the optimization of the tunable text prompts in CoOp.  [21, 24] simultaneously optimize image and text prompts and establish extra connections between different modals.  [22, 47, 3, 53] integrate strong regularization modules or losses into prompt tuning to diminish the overfitting and catastrophic forgetting problem. For better downstream accuracy, researchers design more and more complicated methods, accompanied by inefficiency. To solve the problem, we propose Re-parameterized Low-rank Prompt (RLP) to take the place of classic prompts, which can largely decrease the number of tunable parameters and further enhance the model's generalization ability. Notice that though becoming more complex, existing methods are still refined on a common fundamental basis, i.e. prompt tuning. Such a common basis guarantees that RLP could be easily and smoothly integrated into most of the off-the-shelf methods besides individually applied.
4 Methodology
4.1 A Review of Prompt Tuning for CLIP
In this subsection, we first make a brief review of CLIP [37] which is used as our foundational model. CLIP consists of a text encoder $\mathcal{L}$ and an image encoder $\mathcal{V}$. Typically, $\mathcal{L}$ is a language transformer, while $\mathcal{V}$ can be a convolutional neural network or a vision transformer. In this paper, we follows  [54, 55] to use a ViT-B/16 [7] as the image encoder $\mathcal{V}$ unless specifically mentioned. We will make a review of how to prompt a CLIP for downstream image recognition in the following paragraphs.
Text Encoder Suppose there are $M$ layers in the text encoder. For $k$-th layer $\mathcal{L}_{k}$, the inputs are a series of prompt tokens $P^{l}_{k-1}$ and a [CLS] token $c^{l}_{k-1}$, and the outputs are $P^{l}_{k}$ and $c^{l}_{k}$. The inputs of the first layer $P^{l}_{0}$ and $c^{l}_{0}$ are exactly the word embeddings of the prompts along with the label, e.g. ``A photo of a [CLS]'' or just some randomly initialized vectors. Formally, we have $P^{l}_{k}\in\mathbb{R}^{n^{l}\times d^{l}}$ and $c^{l}_{k}\in\mathbb{R}^{d^{l}}$, where $n^{l}$ denotes the text prompts' length and $d^{l}$ denotes the dimension of word embedding. $\forall 1\leq k\leq M$, we have
$[P^{l}_{k},c^{l}_{k}]=\mathcal{L}_{k}([P^{l}_{k-1},c^{l}_{k-1}])$ (5)
The output feature of the text encoder $f^{l}\in\mathbb{R}^{d^{v}}$, where $d^{v}$ is the dimension of the visual feature space, is generated by projecting the [CLS] token of the last layer to the visual space by a linear transformation, i.e. $f^{l}={\rm Proj}(c^{l}_{M})$.
Image Encoder Suppose there are $N$ layers in the image encoder. For $k$-th layer $\mathcal{V}_{k}$, the inputs are a series of image patch tokens $I_{k-1}$, a classification token $c^{v}_{k-1}$ and prompt tokens $P^{v}_{k-1}$, and the outputs are $I_{k}$, $c^{v}_{k}$ and $P^{v}_{k}$. The inputs of the first layer $I_{0}$ and $c^{v}_{0}$ are exactly the patch embeddings of the input image and the pre-trained class token. $P^{v}_{0}$ is randomly initialized in general. Formally, we have $I_{k}\in\mathbb{R}^{p\times d^{v}}$, $c^{v}_{k}\in\mathbb{R}^{d^{v}}$ and $P^{v}_{k}\in\mathbb{R}^{n^{v}\times d^{v}}$, where $p$ denotes the number of image patches and $d^{v}$ denotes the dimension of visual embedding. $\forall 1\leq k\leq N$, we have
$[P^{v}_{k},c^{v}_{k},I_{k}]=\mathcal{V}_{k}([P^{v}_{k-1},c^{v}_{k-1},I_{k-1}])$ (6)
The output feature of the image encoder is $f^{v}=c^{v}_{N}$.
Prediction CLIP can be easily used for image classification. Suppose there are $C$ classes, and $\{f^{l}_{c}\}_{c=1}^{C}$ are the corresponding text features. The probability of label $y$ is
$p(y|f^{v})=\frac{\mbox{exp}(\mbox{sim}(f^{v},f^{l}_{y})/\tau)}{\sum_{c=1}^{C}\mbox{exp}(\mbox{sim}(f^{v},f^{l}_{c})/\tau)}$ (7)
where sim(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denotes cosine similarity function and $\tau$ is temperature. The final prediction is $\hat{z}=\mathop{\arg\max}\limits_{1\leq y\leq C}(p(y|f^{v}))$.
It is worth noting that some researchers adopt a deeper manner [18, 21] to organize the prompts. They directly add and tune the prompt in each layer in the feature encoder, instead of inheriting the output prompt calculated by the last encoder, i.e. a forward pass becomes
[_$[\_,c^{l}_{k}]=\mathcal{L}_{k}([P^{l}_{k-1},c^{l}_{k-1}])$ (8)
and
[_$[\_,c^{v}_{k},I_{k}]=\mathcal{V}_{k}([P^{v}_{k-1},c^{v}_{k-1},I_{k-1}])$ (9)
Each $P^{l}$/$P^{v}$ contains tunable parameters.
4.2 Re-parameterized Low-rank Prompts
4.2.1 Low-rank Decomposition
Motivated by the observation in Sec. 2, we re-parameterize the origin full-rank prompt matrix into a low-rank form in adaptation for better generalizability and efficiency.
Formally, given a typical text or image prompt matrix $P\in\mathbb{R}^{n\times d}$, we take inspiration from LoRA [16] to create two low-dimentional matrices and use their product as the final low-rank prompt. Suppose the intrinsic rank is $r$, then we first randomly initialize $P_{A}\in\mathbb{R}^{n\times r}$ and $P_{B}\in\mathbb{R}^{r\times d}$. The low-rank prompt $P_{lr}$ is
$P_{lr}=P_{A}P_{B}$ (10)
Here $P_{lr}$ is in the same shape with normal full rank prompt $P$. As a result, we could do a low-rank prompt tuning by sending $P_{lr}$ rather than the original $P$ into the corresponding feature encoder as shown in Eqs. 5 and 6, and updating the parameters in $P_{A}$ and $P_{B}$ through the back-propagated gradients.
4.2.2 Full-rank Initialization
Figure 3: To take full advantage of some special initialization form of prompts, e.g. a hand-crafted template a photo of a for the text prompts, we introduce a concurrent full-rank prompt branch along with the proposed low-rank prompts. By turning off the gradient of the newly added branch, we start training from a promising initial point, and the total number of tunable parameters or stored parameters will not increase as well. The Dropout layer could effectively regularize the update of low-rank prompts and alleviate overfitting and catastrophic forgetting. Better still, Dropout is a lightweight non-parametric layer and turns out to be an Identity layer in inference, resulting in negligible cost.
In the field of tuning vision-language models, existing works have confirmed that the initialization method of prompts is quite important. For example,  [55, 54] adopt a hand-crafted template as the initial point of the text prompts, and  [24] copies the parameters in the text or image class token to initialize the prompts of the corresponding branch. If we just do a random initialization, the overall performance of such existing methods would drop severely. See Sec. 5.2 for more details. In other words, it would be helpful if we could take advantage of a good initial point.
The problem lies in the fact that artificially designed initialization $P_{init}$ is almost certain to be full rank, i.e. $\mbox{rank}(P_{init})=min(n,d)$, and our current low-rank prompt $P_{lr}=P_{A}P_{B}$ could only express the matrices whose rank is less or equal than $r$. Since $r<min(n,d)$, it is impossible to directly initialize $P_{A}$ and $P_{B}$ by a given $P_{init}$.
To solve such a problem, we add a concurrent branch of full-rank prompts $P_{fr}\in\mathbb{R}^{n\times d}$ along with the proposed low-rank prompts as shown in Fig. 3. Naturally, $P_{fr}=P_{init}$. And we randomly sample $P_{A}$/$P_{B}$ from a Gaussian distribution in which $\mu=0$ and $\sigma\rightarrow 0$. A small $\sigma$ here could avoid constant initialization and enrich the update paths. Notably, $P_{fr}$ is kept frozen during the whole adaptation, and thus it would not increase the number of tunable or stored parameters.
4.2.3 Regularization
As discussed in Sec. 3, existing works have shown that proper regularization would significantly improve the generalization ability. Therefore, to alleviate overfitting and catastrophic forgetting, we put a Dropout layer with drop ratio $p$ after the low-rank branch as displayed in Fig. 3.
Therefore, the input prompt of the feature encoder is
$P=P_{fr}+\mbox{Dropout}(P_{lr},p)$ (11)
Finally, we have
$P=P_{init}+\mbox{Dropout}(P_{A}P_{B},p)$ (12)
4.3 Efficiency Analysis
The whole adaptation process of a pre-trained vision-language model can be divided into three parts: training, storage and inference. In this subsection, we will analyze the efficiency of CLP in each part separately.
Training In the training phase, the number of tunable parameters in total is $r(n+d)$. Compared with classic prompt tuning which has $nd$ tunable parameters, we can largely reduce the trainable parameters by choosing a small $r$ satisfying $r<<min(n,d)$. Such reduction could help us train the model faster and use less memory during training than those delicate-designed methods with massive complicated structures and update rules.
Storage After training, we just store $P_{A}$ and $P_{B}$ onto the disk for every downstream task. Similar to the training period, the number of stored parameters is $r(n+d)$. In contrast, classic prompt tuning stores a much larger set of $nd$ parameters. As a result, we can save a fairly large amount of storage space when there are lots of tasks to be adapted.
Inference Before inference, we first load $P_{init}$, $P_{A}$ and $P_{B}$ from disk to memory. Noticing that Dropout is exactly an identity layer in the inference mode, we could pre-calculate the equivalent $P$ by
$P=P_{init}+P_{A}P_{B}$ (13)
and just keep $P$ in the memory. For inference, we directly use $P$ as the input prompts, and thus the inference cost is the same as classic prompt tuning. Some existing methods add complex bridges between the isolated parameters to earn extra improvements, e.g. CoCoOp [55]. There ain’t no such thing as a free lunch. They would face slower speed and huge memory occupation in inference time.
4.4 Advantages of RLP
  • Efficiency and Effectiveness To our knowledge, we are the first one to reach comparable or even better performance with state-of-the-art methods using a fairly small number of parameters, i.e. 516. RLP reaches an excellent trade-off between efficiency and effectiveness.
  • Simplicity Replacing the classic prompt by RLP just needs to modify several lines of code. The design of RLP merely changes the structure of prompts. No special loss nor extra module is introduced.
  • Robustness RLP is relatively capable of anti-disturbance. As in Tab. 3, RLP could maximally reserve its knowledge from domain shift.
  • Plug and Play In a classic transformer model, RLP only affects its prompts, which are very basic and necessary components in modern transformer-based architectures. Such characteristic enables us to plug RLP into most of the existing prompt tuning methods by simply replacing several lines of code.
5 Experiments
Table 2: Comparisons with latest methods in base-to-new generalization. H: harmonic mean [45]. Overall, RLP shows the strongest performance compared with other lastest tuning methods. RLP reaches highest average harmonic mean accuracy and wins on 7 of 11 datasets with far more fewer parameters.
(a) Average over 11 datasets.
#params Base New H
CLIP 0K 69.34 74.22 71.70
CoOp 2.1K 82.69 63.22 71.66
CoCoOp 35.4K 80.47 71.69 75.83
Adapter 525.5K 82.62 70.97 76.35
ProGrad 8.2K 82.79 68.55 75.00
RLP 0.5K 79.86 74.17 76.91
(b) ImageNet.
Base New H
CLIP 72.43 68.14 70.22
CoOp 76.47 67.88 71.92
CoCoOp 75.98 70.43 73.10
Adapter 76.53 66.67 71.26
ProGrad 77.03 68.80 72.68
RLP 76.07 70.87 73.37
(c) Caltech101.
Base New H
CLIP 96.84 94.00 95.40
CoOp 98.00 89.81 93.73
CoCoOp 97.96 93.81 95.84
Adapter 98.20 93.20 95.63
ProGrad 98.50 91.90 95.09
RLP 97.90 95.03 96.45
(d) OxfordPets.
Base New H
CLIP 91.17 97.26 94.12
CoOp 93.67 95.29 94.47
CoCoOp 95.20 97.69 96.43
Adapter 94.40 94.10 94.25
ProGrad 94.40 95.10 94.75
RLP 95.17 97.77 96.45
(e) StanfordCars.
Base New H
CLIP 63.37 74.89 68.65
CoOp 78.12 60.40 68.13
CoCoOp 70.49 73.59 72.01
Adapter 77.13 69.23 72.97
ProGrad 79.00 67.93 73.05
RLP 70.00 75.10 72.46
(f) Flowers102.
Base New H
CLIP 72.08 77.80 74.83
CoOp 97.60 59.67 74.06
CoCoOp 94.87 71.75 81.71
Adapter 97.70 70.83 82.13
ProGrad 96.27 71.07 81.77
RLP 94.80 74.97 83.72
(g) Food101.
Base New H
CLIP 90.10 91.22 90.66
CoOp 88.33 82.26 85.19
CoCoOp 90.70 91.29 90.99
Adapter 90.40 90.40 90.40
ProGrad 90.17 89.53 89.85
RLP 90.73 91.83 91.28
(h) FGVCAircraft.
Base New H
CLIP 27.19 36.29 31.09
CoOp 40.44 22.30 28.75
CoCoOp 33.41 23.71 27.74
Adapter 39.57 32.27 35.55
ProGrad 42.63 26.97 33.04
RLP 35.17 35.57 35.37
(i) SUN397.
Base New H
CLIP 69.36 75.35 72.23
CoOp 80.60 65.89 72.51
CoCoOp 79.74 76.86 78.27
Adapter 81.67 73.93 77.61
ProGrad 80.70 71.03 75.56
RLP 79.47 77.70 78.57
(j) DTD.
Base New H
CLIP 53.24 59.90 56.37
CoOp 79.44 41.18 54.24
CoCoOp 77.01 56.00 64.85
Adapter 80.47 52.23 63.35
ProGrad 76.70 46.67 58.03
RLP 74.27 60.33 66.58
(k) EuroSAT.
Base New H
CLIP 56.48 64.05 60.03
CoOp 92.19 54.74 68.69
CoCoOp 87.49 60.04 71.21
Adapter 86.93 64.20 73.86
ProGrad 91.37 56.53 69.85
RLP 81.57 61.67 70.23
(l) UCF101.
Base New H
CLIP 70.53 77.50 73.85
CoOp 84.69 56.05 67.46
CoCoOp 82.33 73.45 77.64
Adapter 85.80 73.63 79.25
ProGrad 83.90 68.50 75.42
RLP 83.33 75.03 78.97
To verify the effectiveness of the proposed method, we evaluate our method and make comparisons with the latest state-of-the-art methods in terms of the following problems:
  1. 1.
    Base-to-new generalization, which models are trained with base classes and evaluated on both base and new classes.
  2. 2.
    Domain generalization, which aims to show generalization to the domain shift, especially for out-of-distribution data.
  3. 3.
    Cross-dataset transfer, which aims to see if the method has the potential to transfer beyond a single dataset. It is a much more challenging problem because the fundamentals can be totally changed across different datasets.
  4. 4.
    Few-shot learning, which aims to evaluate the adaptation performance of the model to extract knowledge from a dataset whose samples are very few, e.g. 1, 2, 4, 8 or 16 samples.
All methods, including the baseline methods, in our experiments, are initialized with the same CLIP weights provided by the open-source CLIP [37] [^1] https://github.com/openai/CLIP.. In order to give details about the experimental setup, we first introduce the used datasets, the compared baseline methods, and the implementation details.
Datasets
Following previous work [54, 55], we leverage 11 image recognition datasets to verify the effectiveness of the proposed method for both the base-to-new generalization task. These datasets include two datasets for the generic object classification, i.e., ImageNet [5] and Caltech101 [8], five datasets for the fine-grained classification, i.e., OxfordPets [35], StanfordCars [23], Flowers102 [34], Food101 [2] and FGVCAircraft [33], one dataset for the scene recognition, i.e., SUN397 [46], one dataset for the action recognition, i.e., UCF101 [43], one dataset for the texture classification, i.e., DTD [4], and one dataset for the satellite imagery recognition, i.e., EuroSAT [12].
For the domain generalization task, we utilize ImageNet-A [14], ImageNet-R [13], ImageNetv2 [40] and ImageNet-S [44] to verify the robustness of the model. In this setting, we need to first train the model using ImageNet, and then directly use images from other four datasets to do inference.
For the cross-dataset transfer task, the datasets are the same as those of the base-to-new generalization task. Similar to domain generalization, the model will be first trained on ImageNet and then do inference on the other 10 different datasets.
For the few-shot learning task, the datasets are the same as those of the base-to-new generalization task. The model will be trained and evaluated with 1, 2, 4, 8 and 16 shots separately.
The dataset splitting is exactly the same as previous works [54, 55]. We report the averaged model performance over three runs with different random seeds for fair comparisons.
Competitors
  1. 1.
    CLIP [37] (ICML 2021): CLIP is a strong baseline vision-language model that is pre-trained on a large number of image-text pairs from the web by learning a contrastive objective. CLIP enables strong zero-shot adaptation ability on various downstream tasks by using fixed text prompts, i.e. a photo of a.
  2. 2.
    CoOp [54] (IJCV 2022): CoOp replaces the fixed text prompts in CLIP with tunable text prompts to improve the adaptation ability of the vision-language model. CoOp shows excellent performance in few-shot situations.
  3. 3.
    CoCoOp [55] (CVPR 2022): CoCoOp replaces the isolated tunable text prompts in CoOp with conditional text prompts, which receive extra gradients from the image features besides text features. CoCoOp largely improves the generalization ability of the vision-language model, getting good results on base-to-new generalization and domain adaptation.
  4. 4.
    CLIP-Adapter [10] (arxiv): CLIP-Adapter adopts the thoughts of classic Adapter [15] to use serial linear layers and activation functions to adapt for downstream tasks. It is simple yet effective in few-shot learning.
  5. 5.
    ProGrad [56] (ICCV 2023): ProGrad only updates the text and image prompts whose gradient are aligned (or non-conflicting) to the general knowledge, which is represented as the optimization direction offered by the pre-defined prompt predictions. Such regularization helps it finish good adaptation and generalization.
Table 3: Comparisons with latest methods in domain generalization. RLP gets comparable or even better results with the latest state-of-the art methods with much fewer parameters, showing excellent robustness for domain shift.
Source Target
#params ImageNet
-V2
-Sketch
-Adversarial
-Rendition
Average
CLIP 0K 66.73
60.83
46.15
47.77
73.96
57.18
CoOp 2.1K 71.51
64.20
47.99
49.71
75.21
59.28
CoCoOp 35.4K 71.02
64.07
48.75
50.63
76.18
59.91
Adapter 525.5K 69.33
62.53
47.67
49.17
75.42
58.70
ProGrad 8.2K 72.24 64.73
47.61
49.39
74.58
59.07
RLP 0.5K 70.80
63.95
49.07 50.97 77.19 60.30
Table 4: Results in the cross-dataset transfer setting. RLP gives the highest accuracy on 6 of 10 datasets, and slightly outperforms CoCoOp by a small margin on average though the source accuracy of RLP is a bit lower. Such result well demonstrates that RLP could maximally extract general and data-agnostic knowledge from given images.
Source Target
# params
ImageNet
Caltech101
Pets
Cars
Flowers
Food101
Aircraft
Sun397
DTD
EuroSAT
UCF101
Average
CoOp 2.1K 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
CoCoOp 35.4K 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74
Adapter 525.5K 69.33 93.43 88.87 64.40 70.27 85.63 24.67 65.80 44.90 47.70 66.00 65.17
RLP 0.5K 70.57 94.20 90.50 67.17 71.27 86.07 23.83 67.60 46.73 42.10 68.93 65.84
Training Details
Following previous work [55], we employ ViT-B/16 as the image encoder in the CLIP. Each training image is resized to $224\times 224$ before being fed into the image encoder. Some common data augmentation strategies, e.g., random crop and random flip, are used to enhance the model performance, following [55]. During training, we set the batch size as 32. We employ the stochastic gradient descent algorithm (SGD) to optimize the learnable parameters. As [54], we utilize a warm-up scheme at the first epoch, which is important for the tuning of prompts. For all the other baselines, we strictly follow the configurations of their original papers.
For RLP, we conduct a grid search to find the optimal hyper-parameters based on the configuration of CoOp. We chose CoOp, the most lightweight prompt-tuning method, as well as the weakest method as our foundational model. In Sec. 5.1, we set the rank $r=1$ in RLP for all the settings in Sec. 5.1. The total number of tunable parameters in RLP is $4*1+1*512=516$, which is merely around 0.25x of that in CoOp. In Sec. 5.2, we will show how to turn such a weak baseline CoOp into an effective and efficient method step by step, only with the support of such $516$ parameters.
5.1 Main Results
5.1.1 Base-to-new Generalization
Following previous works [55, 54], for each dataset, we split its classes equally into two non-overlapping groups, i.e., one as base classes and the other as new classes. We train all models on the base classes and perform a base/new evaluation on the base/new classes.
The results are shown in Tab. 2. Overall, RLP still shows the strongest performance, and the superiority mainly relies on the improvement of new classes. In other words, RLP largely improves the generalization ability of CLIP. In particular, compared with our direct baseline CoOp, RLP gets $10.95$% accuracy gain on the new classes and $2.83$% accuracy drop on the base classes. Such a result tells us that our low-rank prompt and lightweight regularization design reach our initial purpose to make the model more generalizable, instead of overfitting every upcoming dataset.
Notably, compared with the latest lightweight method ProGrad, even if RLP only uses $<$20x of its parameters, RLP still outperforms it by a clear margin, which clearly demonstrates the excellent effectiveness and efficiency of our proposed method.
Figure 4: Average results of few-shot learning over 11 datasets.
Table 5: Ablation study on base-to-new generalization setting
Low-rank prompts Full-rank init Dropout #params base new H
CoOp - - - 2.1K 82.69 63.22 71.66
RLP - - 0.5K 79.91 71.48 75.46
- 0.5K 79.42 73.21 76.19
0.5K 79.70 73.59 76.53
Table 6: Training, storage, and inference efficiencies.
#params Training throughput Inference throughput
CoOp 2.1K 93 image/s 738 images/s
CoCoOp 35.4K 5 images/s 13 images/s
ProGrad 8.2K 56 images/s 732 images/s
RLP 0.5K 91 images/s 738 images/s
5.1.2 Domain Generalization
Then we follow CoCoOp [55] to use ImageNet, ImageNet-A, ImageNet-R, ImageNet-v2, and ImageNet-S to run domain generalization experiments to verify the robustness of RLP. Results are shown in Tab. 3. As seen from the table, on target datasets, RLP gets comparable or even better average accuracy compared with the latest methods. Importantly, RLP uses much fewer parameters and thus can reach significantly better efficiency. For CLIP, CoOp, CLIP-Adapter and ProGrad, there is a clear performance gap between our RLP and them.
5.1.3 Cross-dataset Transfer
Finally, we follow CoCoOp [55] to conduct cross-dataset transfer evaluation. Results are shown in Tab. 4. We can see that on the source dataset, CoOp accomplishes the highest score, which is similar with the situation in the base-to-new generalization setting. However, concentrating too much on the current dataset will absolutely cause overfitting and catastrophic forgetting problems, and finally lead to a severe drop in the performance on those unseen datasets. In this setting, RLP wins on 6 of 10 datasets and its average accuracy is also slightly better than the best competitor CoCoOp. Such result well demonstrates that RLP could maximally extract general and data-agnostic knowledge from given images compared with other prompt-based methods. Considering the huge difference in the parameter numbers, we could summarize that RLP is still the better choice.
5.1.4 Few-shot Learning
In this paragraph, we will show the experiment results of RLP in the few-shot learning setting. This setting is originated from CoOp. Seen from Fig. 4, RLP consistently outperforms zero-shot CLIP, CoOp, and CLIP-Adapter across all the shot numbers. Such results demonstrate the superiority of RLP in adaptation ability when there are few samples in downstream tasks.
Overall, in base-to-new generalization, domain generalization, cross-dataset transfer and few-shot learning, the proposed method can consistently accomplish state-of-the-art performance while enjoying extremely high parameter efficiency, fruitfully demonstrating the effectiveness and efficiency of the proposed method.
5.2 Analysis
5.2.1 Ablation study
We first investigate the impact of each component in the proposed method RLP. We revise the baseline model, i.e., CoOp [54] which adopts full-rank tunable text prompt, step-by-step.
Our modifications start with transforming the text encoder. We first replace the tunable prompt tokens with our low-rank prompt tokens, denoted as ``Low-rank prompts'', for the CoOp method. After replacing the classic prompts with our low-rank prompts, the average Harmonic accuracy over 11 datasets directly improved from 71.66% to 75.46%. The accuracy on the base classes decreases by 2.78% and on the new classes increases by 8.26%. Although the adaptation ability for base classes slightly drops, the generalization ability for new classes raises quite a lot. This phenomenon proves that our low-rank design is fairly beneficial for the model's generalization ability once again.
Then we integrate the re-parameterized full-rank initialization into CoOp. The harmonic mean accuracy improves by 1.27%. Here we can see the importance of special initialization, even if we have already changed the prompt format from full-rank to low-rank.
Finally, we add a delicate selected lightweight regularization layer, Dropout, to the current module. Dropout helps us alleviate the overfitting problem, and avoid catastrophic forgetting. Therefore, we get a very competitive result, 76.53% Harmonic average accuracy, which is completely comparable to some latest lightweight state-of-the-art methods.
5.2.2 Efficiency Comparison
In this paragraph, we will give a comprehensive analysis of all the training, storage, and inference efficiencies for RLP and several existing methods. Since the parameter scale of CLIP-Adapter [10] is significantly larger than others, we do not contain CLIP-Adapter into comparison.
For a fair comparison, we do all the speed tests on the same GPU. Results are shown in Tab. 6. Our proposed RLP shares nearly the same fastest training and inference speeds with the simplest method CoOp, and more importantly, RLP merely uses 0.5K parameters, which is a lot more storage-efficient than other methods. Specially, compared with classic method CoCoOp, RLP enjoys $>$18x training speed, $>$56x inference speed and $<$70x storage usage. Besides complex structures, the huge gap in the inference speed is partly owing to the huge memory cost of CoCoOp, which forces us to adopt a smaller batch size than other methods for CoCoOp. Moreover, compared with the latest method ProGrad, RLP enjoys $>$1.6x training speed, comparable inference speed, and $<$60x storage usage, which adequately demonstrates the super efficiency of RLP.
Figure 5: Top line: Results of different training epochs on base-to-new generalization. Bottom line: Results of different dropout ratios on base-to-new generalization.
Table 7: Average results on 11 datasets for CLIP with convolutional image encoder, ResNet-50 [11].
image encoder #params base new H
CoOp ResNet-50 2.1K 77.16 61.01 68.14
ProGrad ResNet-50 8.2K 73.29 65.96 69.06
RLP ResNet-50 0.5K 75.22 64.98 69.73
Table 8: Results of adding RLP to image prompts on ImageNet.
#params base new H
Text RLP 0.5K 75.87 70.67 73.17
Image RLP 0.5K 74.57 69.40 71.89
5.2.3 Results for Convolutional Image Encoder
In this subsection, we will show the experiment results of RLP on CLIP with convolutional image encoder ResNet-50 [11], rather than the default ViT-B/16 [7]. Seen from Tab. 7, compared with baseline CoOp, RLP still largely improves the new accuracy and the harmonic mean accuracy over 11 datasets, while the base accuracy slightly drops. Compared with the latest method ProGrad, RLP shows clear superiority on base accuracy and the harmonic mean accuracy.
5.2.4 Results for Using RLP on the Image Side
There are several works [18, 21] indicating the effectiveness of the image prompts. Therefore, in this subsection, we will explore the results of applying RLP to image prompts. Seen from Tab. 8, using RLP on the image side also reaches high accuracy. However, the base, new, and average accuracy of image RLP is not as good as those of text RLP. Such experiment result tells us to use RLP on the text side instead of the image side when we aim to reach high accuracy with extremely few parameters.
5.2.5 Results across Different Hyper-parameters
Effect of Different Epochs
In this paragraph, we investigate that how the total training epoch could influence the adaptation result. As shown in Fig. 5, we select a series of epoch numbers, i.e. 25, 50, 75, 100, and 200. And we run an experiment using RLP to get the corresponding results for the given epoch. We could see that as the training epoch increases, the accuracy on base classes continues decreasing while the accuracy on new classes continues increasing. It is reasonable because as the training continues, the model has a higher and higher risk of forgetting its original knowledge and overfitting the current dataset. To reach a better harmonic mean accuracy, a good trade-off is necessary.
Effect of Different Dropout Ratios
In this paragraph, we will show the influence of different dropout ratios in RLP on ImageNet. Seen from Fig. 5, as the dropout ratio increases, the base accuracy starts decreasing while the new accuracy starts increasing in most cases. The harmonic mean accuracy first increases and then decreases. Such a result is reasonable. Since a high dropout ratio can be regarded as a kind of strong regularization, proper regularization will help avoid overfitting and catastrophic forgetting problems. However, too strong regularization will damage the ability of the model.
Table 9: Results of different ranks on ImageNet.
rank #params base new H
1 0.5K 75.87 70.67 73.17
2 1.0K 76.03 70.13 72.96
3 1.5K 76.20 70.40 73.19
Table 10: Results of different combinations of learning rates and weight decays under 16-shot learning setting on ImageNet
lr wd Acc 1e-4 5e-4 1e-3
1e-3 70.73 70.67 70.78
2e-3 70.80 70.83 70.77
3e-3 70.83 70.83 70.81
Effect of Different Ranks
In this paragraph, we will show the influence of different ranks in RLP. Seen from Tab. 9, roughly, a larger rank brings more parameters, higher base accuracy, and lower new accuracy. As a result, increasing rank is not necessarily able to improve the average accuracy.
Effect of Different Learning Rates and Weight Decays
In this paragraph, we investigate the effect of normal hyper-parameters learning rate and weight decay. We wonder if tuning them carefully would lead to significant improvement. Seen from Tab. 10, the performance of RLP stays stable whatever the learning rate and weight decay vary. The conclusion here is that our method RLP is robust for learning rate and weight decay.
6 Conclusion
With the development of huge vision-language models, how to effectively and efficiently adapt such huge models to various downstream tasks becomes a challenging problem. Much effort has been made to leverage the potential of prompt tuning in adapting vision-language models. However, existing methods usually suffer from inefficiency. To solve such a problem, we propose Re-parameterized Low-rank Prompts (RLP) for a better trade-off between effectiveness and efficiency in vision-language adaptation. Along with the low-rank prompt design, we further propose a novel initialization method and integrate the Dropout layer as a lightweight regularization module to further improve the performance without introducing any extra parameters and inference cost. We are the first to explore how to effectively adapt vision-language models using an extremely small number of parameters, i.e. 0.5K. Besides efficiency and effectiveness, RLP has many other advantages such as simplicity and robustness, which are quite valuable in real-world application scenarios. Extensive experiments show that even if RLP uses much fewer parameters, RLP could still consistently be comparable to or better than state-of-the-art methods.